Circular No. 7666

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA CYGNI 2001 H. Sato, National Astronomical Observatory (NAO), Tokyo, reports

the discovery of a probable nova (mag 11.9) by Yuji Nakamura (Mie, Japan) on a Tri-X film exposed on July 13.651 UT with a 200-mm f/4 lens. The object was also present on CCD exposures taken with a 0.5-m reflector by H. Fukushima, NAO, resulting in the following magnitudes: July 16.515, V = 11.69; 16.533, I = 9.55. H. Yamaoka, Kyushu University, forwards the following precise position for the variable by K. Kadota from a CCD image taken on July 17.735: α = 20^{h}07^{m}17^{s}.94, δ = +36^{o}04′37′′.2 (equinox 2000.0; twelve GSC-ACT reference stars). Yamaoka adds that no distinct object is visible on the Digital Sky Survey within 3′′ of this position.

T. Iijima, Osservatorio Astronomico di Padova, writes: “Optical spectra (range 430–680 nm) were taken on July 17.9 UT using a Boller and Chivens spectrograph mounted on the 1.22-m telescope of the Osservatorio Astrofisico di Asiago. The spectra show broad emission lines of H I, Fe II, and probably Na I D and N II. Some prominent emission lines are accompanied by P-Cyg-type absorption components. The FWHM of the Hα emission component is ∼ 950 km/s, and a weak absorption component blueshifted by ∼ −1200 km/s is seen. This object could be a classical nova in the early decline stage from maximum light.”

SUPERNOVA 2001cy IN UGC 11927 N. V. Borisov, V. V. Vlasuk, A. N. Burenkov, and D. K. Klochkov,

Special Astrophysical Observatory, Russian Academy of Sciences, report that a spectrum (range 360–810 nm, dispersion 0.24 nm/pixel, S/N = 70; 15-min exposure) of SN 2001cy (cf. IAUC 7655, 7665), obtained on July 13.03 UT with the 6-m telescope (+ long slit spectrograph + CCD camera), shows it to be a type-II supernova. Hα emission from the supernova indicates a recession velocity of 1900 km/s for the host galaxy.

COMET C/2001 A2 (LINEAR) L. M. Woodney and D. G. Schleicher, Lowell Observatory; and R.

Greer, Wittenberg University, report narrowband gas and dust imaging of this comet: “On June 29–30, the comet displayed CN jet(s) symmetrical about p.a. 250^{o}. Three successive arcs separated by ≈ 12 000 km were observed on each side; outward motion of the arcs was detected. These arcs were not observed in the dust continuum.”

2001 July 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green